OpenStudy (anonymous):

Solve the equation, Check for extraneous solutions. 9| 9-8x| = 2x+3

OpenStudy (anonymous):

Please help don't understand!

OpenStudy (debbieg):

Isolate the abs. value part, first. You ALWAYS need an equation of the form: |something|= something

OpenStudy (anonymous):

how do you solve with the 9 there?

OpenStudy (debbieg):

Sorry, my browser crashed or something and I couldn't respond. I told you, you need to isolate the abs value first. So how can you get the 9 "away" from the left hand side? what must you do to both sides?

OpenStudy (debbieg):

Remember that the absolute value signs are really a "grouping symbol".... so you have \(\Large 9\cdot| 9-8x| = 2x+3 \) Now, in general, if you have \(\Large 9\cdot A = B \) and you want to "isolate A", what do you do?